# Eugene Lerman

### v1

@@ 1. Introduction
In the case of singular symplectic quotients the approaches roughly fall into three categories: “geometric” “algebraic” (cf. [AGJ]) and “homological” [St]. In this paper we connect geometric and algebraic approaches using tools of algebraic geometry over $C^{\infty}$-rings [Du, J]. (The connections with homological approach should come out of derived differential geometry once it is well developed.) From the $C^{\infty}$-ring point of view the algebraic approach results in Poisson $C^{\infty}$-rings (a concept that we introduce in this paper) while geometric approach results in spaces with a structure sheaf of Poisson $C^{\infty}$-rings. We show that Dubuc’s spectrum functor from the category of $C^{\infty}$-rings to the category of local $C^{\infty}$-ringed spaces send Poisson $C^{\infty}$-rings to Poisson ringed spaces (see Theorem 3.18). This unifies algebraic and geometric views of reduction in symplectic and Poisson geometry. It also makes singular reduction a lot simpler since there is no need to worry about stratifications. With almost no extra work we also makes sense of singular quasi-Poisson reduction as Poisson $C^{\infty}$-schemes. Other applications include coisotropic reduction in Poisson geometry and Poisson-Dirac submanifolds and more generally subschemes.
@@
Recall that a Poisson algebra $A$ over a commutative ring $k$ is an associative (and often commutative) $k$-algebra with a compatible Lie algebra bracket {,}$\{\cdot,\cdot\}:A\times A\to A$ such that for each $a\in A$ the map $\{a,\cdot\}:A\to A$ is a $k$-algebra derivation. Its is well-known [Po, K1] that the algebra-geometric spectrum of a Poisson algebra is a Poisson scheme, that is, a local ringed space whose structure sheaf is a sheaf of Poisson algebras. Poisson schemes have proven to be useful in algebraic geometry and representation theory. However, if one starts with a Poisson manifold $M$ and regards its algebra of functions $C^{\infty}(M)$ as Poisson algebra over $\mathbb{R}$ then the associated Poisson scheme is not the original manifold with its sheaf of smooth functions. Of course one may feel that there is no need to view Poisson manifolds as affine Poisson schemes, and it’s hard to argue with this point of view. On the other hand the Sniatycki-Weinstein algebraic reduction [SW] produces a Poisson algebra over $\mathbb{R}$, which is a Poisson algebra of a manifold under a regularity assumption. However, in general this Poisson algebra need not be an algebra of functions on any space [AGJ]. It is natural to wonder if this algebra is the algebra of global sections of a sheaf on some space. This question was the original motivation for undertaking the work in this paper.
@@
There are two appendices. In Appendix A we address the question of getting one’s hands on $\operatorname{Spec}(\mathscr{A})$. We show that if a $C^{\infty}$-ring $\mathscr{A}$ is finitely generated (as a $C^{\infty}$-ring!) then the space $X_{\mathscr{A}}$ of its points embeds as a closed subspace in some $\mathbb{R}^{N}$ (the embedding depends on the choice of generators). If $G\times M\to M$ is an action of a Lie group on a manifold $M$ then the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of invariant functions may or may not be finitely generated (if $G$ is compact and the number of orbit types is finite then $C^{\infty}(M)^{G}$ is finitely generated by a nontrivial theorem of Schwarz [Sch]). Regardless of $C^{\infty}(M)^{G}$ being finitely generated we show that for proper actions (of not necessarily compact) Lie groups the space $X_{C^{\infty}(M)^{G}}$ of points the ring $C^{\infty}(M)^{G}$ of invariants is homeomorphic to the orbit space $M/G$ and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})$ is isomorphic to $(M/G,C^{\infty}_{M/G})$ where $C^{\infty}_{M/G}\subset C^{0}_{M/G}$ is a sheaf of “smooth” functions on the orbit space. This implies, in particular, if $M$ is Poisson, the action of $G$ on $M$ is Hamiltonian, $\mu:M\to\mathfrak{g}^{\vee}$ is an equivariant moment map, $\alpha\in\mathfrak{g}^{\vee}$ and $I_{\alpha}\subset C^{\infty}(M)^{G}$ is the vanishing ideal of the level set $\mu^{-1}(\alpha)$ then the set of points $X_{C^{\infty}(M)^{G}/I_{\alpha}}$ embeds into the orbit space $M/G$ as a closed subspace and the $C^{\infty}$-Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/I_{\alpha})$ “is” a closed subscheme of $(M/G,C^{\infty}_{M/G})$. The same result, mutatis mutandis holds if $M$ is a quasi-Poisson manifold.
In Appendix B we discuss the Cushman-Weinstein conjecture [Eg1, Eg2, SL] on embedding of symplectic quotients into Euclidean spaces.
@@ 2. Poisson $C^{\infty}$-rings
We start by introducing notation, which is not entirely standard. It is, however, convenient.
@@
[…]
@@
Another application of Lemma 2.14 gives us the so called universal reduction of Arms, Cushman and Gotay [ACG]:
@@
the set of all invariant functions that vanish on the level set $\mu^{-1}(\alpha)$. Since the moment map $\mu$ is constant along the integral curves of all invariant functions $f\in C^{\infty}(M)^{G}$, the flows of invariant functions preserve $\mu^{-1}(\alpha)$. By Lemma 2.14 the ideal $I_{\alpha}$ is a Poisson ideal in the Poisson $C^{\infty}$-ring $cin(M)^{G}$. Therefore the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ is a Poisson $C^{\infty}$-ring.
@@
Any element of the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ gives rise to a well-defined continuous function on the quotient space $\mu^{-1}(\alpha)/G_{\alpha}$ where $G_{\alpha}$ is the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ (under the coadjoint action). In the next section (Theorem 3.18) we will see that one can associate to any Poisson $C^{\infty}$-ring $\mathscr{A}$ a space $X_{\mathscr{A}}$ with a sheaf of Poisson $C^{\infty}$-rings $\mathscr{O}_{\mathscr{A}}$. If the action of $G$ on $M$ is nice enough, the two spaces turn out to be homeomorphic. We will prove it in Appendix A.
@@
[…]
@@ 3. Poisson $C^{\infty}$-schemes
The surjectivity of (3.1) follows from the fact that $C^{\infty}(M)$ contains a proper function combined with Lemma A.7. The injectivity of (3.1) is the fact that smooth functions on Hausdorff manifolds separate points.
@@
    for any map $\varphi:\mathscr{A}\to\mathscr{B}$ of $C^{\infty}$-rings so that $\varphi(a)$ is invertible in $\mathscr{B}$ for all $a\in\Sigma$ there exists a unique map $\overline{\varphi}:\mathscr{A}\{\Sigma^{-1}\}\to\mathscr{B}$ making the diagram
    ${\mathscr{A}}$${\mathscr{B}}$${\mathscr{A}\{\Sigma^{-1}\}}$$\scriptstyle{\varphi}$$\scriptstyle{\gamma}$$\scriptstyle{\overline{\varphi}}$
@@
(3.3) $I_{x}:=\{a\in\mathscr{A}\mid\textrm{ there is }d\in\mathscr{A}\textrm{ so that }x(d)\not=0\textrm{ and }ad=0\}.$
@@
(3.5) $\begin{split}\mathscr{O}_{\mathscr{A}}(U)=&\{s:U\to\bigsqcup_{x\in U}\mathscr{A}_{x}\mid\textrm{ there is an open cover }\{U_{\alpha}\}_{\alpha\in A}\textrm{ of }U\\ &\textrm{ and }\{a_{\alpha}\}_{\alpha\in A}\subset\mathscr{A}\textrm{ so that }s|_{U_{\alpha}}=\mathfrak{s}_{a_{\alpha}}|_{U_{\alpha}}\textrm{ for all }\alpha\in A\}.\end{split}$
@@
(3.7) ${\mathscr{A}}$${\mathscr{B}}$${\mathscr{A}_{x\circ\varphi}}$${\mathscr{B}_{x}}$$\scriptstyle{\pi_{x\circ\varphi}}$$\scriptstyle{\pi_{x}}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi_{x}}$
@@ 4. Applications
4.i. Symplectic reduction
@@
In the case where the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is sufficiently nice and the action of $G$ on $M$ is proper the space $X_{\mathscr{A}/I_{\alpha}}$ of points of the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ is homeomorphic to $\mu^{-1}(\alpha)/G_{\alpha}$ (where $G_{\alpha}$ denotes the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ under the coadjoint actions). We’ll discuss it in the next section, Section A. Therefore we may regard the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ as (a possibly singular) symplectic quotient (a.k.a. reduced space) $M/\!/_{\alpha}G$.
@@
for all $t\in\mathbb{R}$ and all $[x,y]\in\mathbb{R}^{2}/\mathbb{Z}^{2}$. Then the lifted action of $\mathbb{R}$ on the cotangent bundle $T^{*}\mathbb{T}^{2}\simeq\mathbb{T}^{2}\times\mathbb{R}^{2}$ is Hamiltonian with the moment $\mu([x,y],p_{x},p_{y})=p_{x}+\sqrt{2}p_{y}$. Consequently
@@
Since $\varpi:C^{\infty}(M)^{G}\to C^{\infty}(M)^{G}/{\mathcal{I}}(X)$ is surjective, Lemma A.5 implies that the induced map $f(\varphi):X_{C^{\infty}(M)^{G}/{\mathcal{I}}(X)}\to X_{C^{\infty}(M)^{G}}$ of spaces is a closed embedding. Furthermore, by Lemma A.10, if the action of $G$ on $M$ is proper then the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to $M/G$. Hence in the case of proper actions the space of points of the affine Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/{\mathcal{I}}(X))$ is homeomorphic to a closed subset of the orbit space $M/G$.
@@
Case a. The ideal $I$ is a Poisson subalgebra of $C^{\infty}(P)$, i.e, for all $f,g\in I$ the bracket $\{f,g\}\in I$. Then the ideal $I$ is contained in its normalizer $N(I)$. Hence the quotient $N(I)/I$ is a Poisson $C^{\infty}$-ring and its spectrum $\operatorname{Spec}(N(I)/I$ is a Poisson $C^{\infty}$-scheme. This is a version of coisotropic reduction in Poisson geometry.
@@
However, suppose $C\subset P$ is a coisotropic submanifold and let $I$ denote the ideal of functions that vanish on $C$. Then $I$ is a Poisson subalgebra of $C^{\infty}(P)$ and the Poisson $C^{\infty}$-ring $N(I)/I$ is isomorphic to the $C^{\infty}$-ring of functions on $C$ that are constant along the leaves of the null foliation $\mathscr{F}$. If $C/\mathscr{F}$ is a manifold then it is isomorphic to $\operatorname{Spec}(N(I)/I)$. In general we view the Poisson $C^{\infty}$-scheme $\operatorname{Spec}(N(I)/I)$ as the coisotropic reduction of $C$. Note that the underlying topological space $X_{N(I)/I}$ may be nicer than the leaf space $C/\mathscr{F}$ since (as was mentioned above) the space of points of a $C^{\infty}$-ring is Hausdorff and completely regular. We have seen an instance of it when we looked at the reduction of the cotangent bundle of the 2-torus $T^{*}\mathbb{T}$ by the irrational flow (Example 4.1).
@@
We now revisit Example 2.21. Recall that one starts with a symplectic manifold $(M,\omega)$ equipped with a Hamiltonian action of a connected Lie group $G$ and associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Then the ideal $\mathscr{I}\subset C^{\infty}(M)$ generated by the components of the moment map $\mu$ is closed under the Poisson brackets, hence $N(\mathscr{I})/\mathscr{I}$ is Poisson $C^{\infty}$-ring equipped with an injective map $\varphi:N(\mathscr{I})/\mathscr{I}\to C^{\infty}(M)/\mathscr{I}$. Hence $\operatorname{Spec}(N(\mathscr{I})/\mathscr{I}$ is a Poisson scheme which we may view as a coisotropic reduction of $C^{\infty}(M)/\mathscr{I}$.
Note that since the quotient map $q:C^{\infty}(M)\to C^{\infty}(M)/\mathscr{I}$ is surjective, the induced map $f(q):X_{C^{\infty}(M)/\mathscr{I}}\to X_{C^{\infty}(M)}\simeq M$ is a closed embedding by Lemma A.5. However, in general the ideal $\mathscr{I}$ may be strictly smaller than the vanishing ideal of $\mu^{-1}(0)$ (cf. [AGJ]) and the scheme $\operatorname{Spec}(C^{\infty}(M)/\mathscr{I})$ need not be reduced.
Case b. The ideal $I$ is an arbitrary ideal of the $C^{\infty}$-ring $C^{\infty}(P)$. Then the intersection $N(I)\cap I\subseteq N(I)$ is a Poisson ideal in $N(I)$. Consequently $\operatorname{Spec}(N(I)/I\cap N(I))$ is a Poisson scheme. Note that the inclusion $N(I)\hookrightarrow C^{\infty}(P)$ induces an injective map
(4.1) $\varphi:N(I)/I\cap N(I)\hookrightarrow C^{\infty}(P)/I.$
If the map $\varphi$ above is an isomorphism then since $N(I)/I\cap N(I)$ is Poisson $C^{\infty}$-ring, the quotient $C^{\infty}(P)/I$ is a also Poisson $C^{\infty}$-ring. Then the close subscheme $\operatorname{Spec}(C^{\infty}(P)/I)$ of the manifold $P$ is a Poisson scheme. When $I$ is a vanishing ideal of a submanifold $M\subset P$, then $\operatorname{Spec}(C^{\infty}(P)/I)=M$ and $M$ is a Poisson-Dirac submanifold of $P$.
@@
Appendix A Computing $\operatorname{Spec}(\mathscr{A})$
We start by recalling what it means for a $C^{\infty}$-ring $\mathscr{A}$ to be finitely generated. For such rings Dubuc [Du] gave a fairly concrete description of $\operatorname{Spec}(\mathscr{A})$ which we describe. Next we consider point determined $C^{\infty}$-rings (see Definition A.2). We show that if $\mathscr{A}$ is point determined then the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of $\operatorname{Spec}(\mathscr{A})$ is a sheaf of functions on the space $X_{\mathscr{A}}$ of its points. For finitely generated and point determined $C^{\infty}$-rings the description of $\operatorname{Spec}(\mathscr{A})$ simplifies considerably: $\operatorname{Spec}(\mathscr{A})$ is a closed subset of some $\mathbb{R}^{N}$ with the induced sheaf of “smooth” functions.

### v2

@@ 1. Introduction
In the case of singular symplectic quotients the approaches roughly fall into three categories: “geometric” “algebraic” (cf. [AGJ]) and “homological” [St]. In this paper we connect geometric and algebraic approaches using tools of algebraic geometry over $C^{\infty}$-rings [Du, J]. (The connections with homological approach should come out of derived differential geometry once it is well developed.) From the $C^{\infty}$-ring point of view the algebraic approach results in Poisson $C^{\infty}$-rings (a concept that we introduce in this paper) while geometric approach results in spaces with a structure sheaf of Poisson $C^{\infty}$-rings. We show that Dubuc’s spectrum functor from the category of $C^{\infty}$-rings to the category of local $C^{\infty}$-ringed spaces send Poisson $C^{\infty}$-rings to Poisson ringed spaces (see Theorem 3.18). This unifies algebraic and geometric views of reduction in symplectic and Poisson geometry. It also makes singular reduction a lot simpler since there is no need to worry about stratifications. With almost no extra work we also makes sense of singular quasi-Poisson reduction as Poisson $C^{\infty}$-schemes. Other applications include coisotropic reduction in Poisson geometry and Poisson-Dirac submanifolds and, more generally, subschemes.
@@
Recall that a Poisson algebra $A$ over a commutative ring $k$ is an associative (and often commutative) $k$-algebra with a compatible Lie algebra bracket $\{\cdot,\cdot\}:A\times A\to A$ such that for each $a\in A$ the map $\{a,\cdot\}:A\to A$ is a $k$-algebra derivation. Its is well-known [Po, K1] that the algebra-geometric spectrum of a Poisson algebra is a Poisson scheme, that is, a local ringed space whose structure sheaf is a sheaf of Poisson algebras. Poisson schemes have proven to be useful in algebraic geometry and representation theory. However, if one starts with a Poisson manifold $M$ and regards its algebra of functions $C^{\infty}(M)$ as Poisson algebra over $\mathbb{R}$ then the associated Poisson scheme is not the original manifold with its sheaf of smooth functions. Of course one may feel that there is no need to view Poisson manifolds as affine Poisson schemes, and it’s hard to argue with this point of view. On the other hand the Sniatycki-Weinstein algebraic reduction [SW] produces a Poisson algebra over $\mathbb{R}$, which is a Poisson algebra of a manifold under a regularity assumption. However, in general this Poisson algebra need not be an algebra of functions on any space [AGJ]. It is natural to wonder if this algebra is the algebra of global sections of a sheaf on some space. This question was the original motivation for undertaking the work in this paper.
@@
In Section 5 we address the question of getting one’s hands on $\operatorname{Spec}(\mathscr{A})$. We show that if a $C^{\infty}$-ring $\mathscr{A}$ is finitely generated (as a $C^{\infty}$-ring!) then the space $X_{\mathscr{A}}$ of its points embeds as a closed subspace in some $\mathbb{R}^{N}$ (the embedding depends on the choice of generators). If $G\times M\to M$ is an action of a Lie group on a manifold $M$ then the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of invariant functions may or may not be finitely generated (if $G$ is compact and the number of orbit types is finite then $C^{\infty}(M)^{G}$ is finitely generated by a nontrivial theorem of Schwarz [Sch]). Regardless of $C^{\infty}(M)^{G}$ being finitely generated we show that for proper actions (of not necessarily compact) Lie groups the space $X_{C^{\infty}(M)^{G}}$ of points the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ of invariants is homeomorphic to the orbit space $M/G$ and consequently $\operatorname{Spec}(C^{\infty}(M)^{G})$ is isomorphic to $(M/G,C^{\infty}_{M/G})$ where $C^{\infty}_{M/G}\subset C^{0}_{M/G}$ is a sheaf of “smooth” functions on the orbit space. This implies, in particular, if $M$ is Poisson, the action of $G$ on $M$ is Hamiltonian, $\mu:M\to\mathfrak{g}^{\vee}$ is an equivariant moment map, $\alpha\in\mathfrak{g}^{\vee}$ and $I_{\alpha}\subset C^{\infty}(M)^{G}$ is the vanishing ideal of the level set $\mu^{-1}(\alpha)$ then the set of points $X_{C^{\infty}(M)^{G}/I_{\alpha}}$ embeds into the orbit space $M/G$ as a closed subspace and the $C^{\infty}$-Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/I_{\alpha})$ “is” a closed subscheme of $(M/G,C^{\infty}_{M/G})$. The same result, mutatis mutandis holds if $M$ is a quasi-Poisson manifold.
In Section 6 we discuss the Cushman-Weinstein conjecture [Eg1, Eg2, SL] on embedding of symplectic quotients into Euclidean spaces as “Poisson subvarieties.”
@@ 2. Poisson $C^{\infty}$-rings
In the introduction we sketched a definition of a $C^{\infty}$-ring. We omitted the definition of $C^{\infty}$-ring morphisms and of subrings of $C^{\infty}$-rings. See [MR, J] for a more careful discussion. In place of such a discussion we introduce some notation, which is not entirely standard. It is, however, convenient.
@@
A Poisson $C^{\infty}$-ring is a $C^{\infty}$-ring $\mathscr{A}$ together with a Poisson bracket $\{\cdot,\cdot\}:\mathscr{A}\times\mathscr{A}\to\mathscr{A}$ on the underlying $\mathbb{R}$-algebra so that for any $a\in\mathscr{A}$ the map
@@
Another application of Lemma 2.14 gives us the so called “universal reduction” of Arms, Cushman and Gotay [ACG]:
@@
the set of all invariant functions that vanish on the level set $\mu^{-1}(\alpha)$. Since the moment map $\mu$ is constant along the integral curves of all invariant functions $f\in C^{\infty}(M)^{G}$, the flows of invariant functions preserve $\mu^{-1}(\alpha)$. By Lemma 2.14 the ideal $I_{\alpha}$ is a Poisson ideal in the Poisson $C^{\infty}$-ring $C^{\infty}(M)^{G}$. Therefore the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ is a Poisson $C^{\infty}$-ring.
@@
Any element of the quotient $C^{\infty}$-ring $C^{\infty}(M)^{G}/I_{\alpha}$ gives rise to a well-defined continuous function on the quotient space $\mu^{-1}(\alpha)/G_{\alpha}$ where $G_{\alpha}$ is the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ (under the coadjoint action). In the next section (Theorem 3.18) we will see that one can associate to any Poisson $C^{\infty}$-ring $\mathscr{A}$ a space $X_{\mathscr{A}}$ with a sheaf of Poisson $C^{\infty}$-rings $\mathscr{O}_{\mathscr{A}}$. If the action of a Lie group $G$ on a manifold $M$ is nice enough, the two spaces turn out to be homeomorphic. We will prove it in Section 5.
@@
Let $(M,\omega)$ be a symplectic manifold with a Hamiltonian action of a connected Lie group $G$ and associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Denote the canonical pairing $\mathfrak{g}^{\vee}\times\mathfrak{g}\to\mathbb{R}$ by $\langle\cdot,\cdot\rangle$. For every element $X$ in the Lie algebra $\mathfrak{g}$ we have a smooth function $\mu_{X}:=\langle\mu,X\rangle:M\to\mathbb{R}$. Let $\mathscr{I}$ be the ideal in the $C^{\infty}$-ring $C^{\infty}(M)$ generated by the set $\{\mu_{X}\mid X\in\mathfrak{g}\}$. Since $\mu$ is equivariant, $\{\mu_{X},\mu_{Y}\}=\mu_{[X,Y]}$ for all $X,Y\in\mathfrak{g}$. Consequently $\mathscr{I}$ is a Poisson subalgebra of $C^{\infty}(M)$. Sniatycki and Weinstein define the reduced algebra associated to the data $(M,\omega,\mu)$ to be the algebra
@@ 3. Poisson $C^{\infty}$-schemes
The surjectivity of (3.1) follows from the fact that $C^{\infty}(M)$ contains a proper function combined with Lemma 5.7. The injectivity of (3.1) is the fact that smooth functions on Hausdorff manifolds separate points.
@@
    for any map $\varphi:\mathscr{A}\to\mathscr{B}$ of $C^{\infty}$-rings so that $\varphi(a)$ is invertible in $\mathscr{B}$ for all $a\in\Sigma$ there exists a unique map $\overline{\varphi}:\mathscr{A}\{\Sigma^{-1}\}\to\mathscr{B}$ of $C^{\infty}$-rings making the diagram
    ${\mathscr{A}}\textstyle{\mathscr{B}}$${\mathscr{A}\{\Sigma^{-1}\}}$$\scriptstyle{\varphi}$$\scriptstyle{\gamma}$$\scriptstyle{\overline{\varphi}}$
@@
[…]
@@
[…]
@@
(3.7) ${\mathscr{A}}\textstyle{\mathscr{B}}$${\mathscr{A}_{x\circ\varphi}}$${\mathscr{B}_{x}}$$\scriptstyle{\pi_{x\circ\varphi}}$$\scriptstyle{\pi_{x}}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi_{x}}$
@@ 4. Applications
4.i. Marsden-Weinstein-Meyer reduction
@@
In the case where the $C^{\infty}$-ring $C^{\infty}(M)^{G}$ is sufficiently nice and the action of $G$ on $M$ is proper the space $X_{\mathscr{A}/I_{\alpha}}$ of points of the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ is homeomorphic to $\mu^{-1}(\alpha)/G_{\alpha}$ (where $G_{\alpha}$ denotes the stabilizer of $\alpha\in\mathfrak{g}^{\vee}$ under the coadjoint actions). We’ll discuss it in the next section, Section 5. Therefore we may regard the scheme $\operatorname{Spec}(\mathscr{A}/I_{\alpha})$ as (a possibly singular) symplectic quotient (a.k.a. reduced space) $M/\!/_{\alpha}G$.
@@
for all $t\in\mathbb{R}$ and all $x,y]\in\mathbb{R}^{2}/\mathbb{Z}^{2}$. Then the lifted action of $\mathbb{R}$ on the cotangent bundle $T^{*}\mathbb{T}^{2}\simeq\mathbb{T}^{2}\times\mathbb{R}^{2}$ is Hamiltonian with the moment $\mu([x,y],p_{x},p_{y})=p_{x}+\sqrt{2}p_{y}$. Consequently
@@
Since $\varpi:C^{\infty}(M)^{G}\to C^{\infty}(M)^{G}/{\mathcal{I}}(X)$ is surjective, Lemma 5.5 implies that the induced map $f(\varphi):X_{C^{\infty}(M)^{G}/{\mathcal{I}}(X)}\to X_{C^{\infty}(M)^{G}}$ of spaces is a closed embedding. Furthermore, by Lemma 5.10, if the action of $G$ on $M$ is proper then the space of points $X_{C^{\infty}(M)^{G}}$ is homeomorphic to $M/G$. Hence in the case of proper actions the space of points of the affine Poisson scheme $\operatorname{Spec}(C^{\infty}(M)^{G}/{\mathcal{I}}(X))$ is homeomorphic to a closed subset of the orbit space $M/G$.
@@
Case a. The ideal $I$ is a Poisson subalgebra of $C^{\infty}(P)$, i.e, for all $f,g\in I$ the bracket $\{f,g\}\in I$. Then the ideal $I$ is contained in its normalizer $N(I)$. Hence the quotient $N(I)/I$ is a Poisson $C^{\infty}$-ring and its spectrum $\operatorname{Spec}(N(I)/I)$ is a Poisson $C^{\infty}$-scheme. This is a version of coisotropic reduction in Poisson geometry.
@@
However, suppose $C\subset P$ is a coisotropic submanifold and let $I$ denote the ideal of functions that vanish on $C$. Then $I$ is a Poisson subalgebra of $C^{\infty}(P)$ and the Poisson $C^{\infty}$-ring $N(I)/I$ is isomorphic to the $C^{\infty}$-ring of functions on $C$ that are constant along the leaves of the null foliation $\mathscr{F}$. If the space of leaves $C/\mathscr{F}$ is a manifold then $C/\mathscr{F})$ is isomorphic to $\operatorname{Spec}(N(I)/I)$. In general we view the Poisson $C^{\infty}$-scheme $\operatorname{Spec}(N(I)/I)$ as the coisotropic reduction of $C$. Note that the underlying topological space $X_{N(I)/I}$ may be nicer than the leaf space $C/\mathscr{F}$ since (as was mentioned above) the space of points of a $C^{\infty}$-ring is Hausdorff and completely regular. We have seen an instance of it when we looked at the reduction of the cotangent bundle of the 2-torus $T^{*}\mathbb{T}$ by the irrational flow (Example 4.1).
@@
We now revisit Example 2.21. Recall that one starts with a symplectic manifold $(M,\omega)$ equipped with a Hamiltonian action of a connected Lie group $G$ and associated equivariant moment map $\mu:M\to\mathfrak{g}^{\vee}$. Then the ideal $\mathscr{I}\subset C^{\infty}(M)$ generated by the components of the moment map $\mu$ is closed under the Poisson brackets, hence $N(\mathscr{I})/\mathscr{I})$ is Poisson $C^{\infty}$-ring equipped with an injective map $\varphi:N(\mathscr{I})/\mathscr{I}\to C^{\infty}(M)/\mathscr{I}$. Hence $\operatorname{Spec}(N(\mathscr{I})/\mathscr{I}$ is a Poisson scheme which we may view as a coisotropic reduction of $C^{\infty}(M)/\mathscr{I}$.
Note that since the quotient map $q:C^{\infty}(M)\to C^{\infty}(M)/\mathscr{I}$ is surjective, the induced map $f(q):X_{C^{\infty}(M)/\mathscr{I}}\to X_{C^{\infty}(M)}\simeq M$ is a closed embedding by Lemma 5.5. However, in general the ideal $\mathscr{I}$ may be strictly smaller than the vanishing ideal of $\mu^{-1}(0)$ (cf. [AGJ]) and the scheme $\operatorname{Spec}(C^{\infty}(M)/\mathscr{I})$ need not be reduced.
Case b. The ideal $I$ is an arbitrary ideal of the $C^{\infty}$-ring $C^{\infty}(P)$. Then the intersection $N(I)\cap I\subseteq N(I)$ is a Poisson ideal in $N(I)$. Consequently $\operatorname{Spec}(N(I)/\left(I\cap N(I))\right)$ is a Poisson scheme. Note that the inclusion $N(I)\hookrightarrow C^{\infty}(P)$ induces an injective map
(4.1) $\varphi:N(I)/(I\cap N(I))\hookrightarrow C^{\infty}(P)/I.$
If the map $\varphi$ above is an isomorphism then, since $N(I)/(I\cap N(I))$ is Poisson $C^{\infty}$-ring, the quotient $C^{\infty}(P)/I$ is a also Poisson $C^{\infty}$-ring. Consequently the close subscheme $\operatorname{Spec}(C^{\infty}(P)/I)$ of the manifold $P$ is a Poisson scheme. When $I$ is a vanishing ideal of a submanifold $M\subset P$, then $\operatorname{Spec}(C^{\infty}(P)/I)=M$ and $M$ is a Poisson-Dirac submanifold of $P$.
@@
5. Computing $\operatorname{Spec}(\mathscr{A})$
Here is a brief summary of the content of this section We start by recalling what it means for a $C^{\infty}$-ring $\mathscr{A}$ to be finitely generated. For such rings Dubuc [Du] gave a fairly concrete description of $\operatorname{Spec}(\mathscr{A})$ which we recall. Next we consider point determined $C^{\infty}$-rings (see Definition 5.2). We show that if $\mathscr{A}$ is point determined then the structure sheaf $\mathscr{O}_{\mathscr{A}}$ of $\operatorname{Spec}(\mathscr{A})$ is a sheaf of functions on the space $X_{\mathscr{A}}$ of its points. For finitely generated and point determined $C^{\infty}$-rings the description of $\operatorname{Spec}(\mathscr{A})$ simplifies considerably: $\operatorname{Spec}(\mathscr{A})$ is a closed subset of some $\mathbb{R}^{N}$ with the induced sheaf of “smooth” functions.